# UniversalCEFR: Enabling Open Multilingual Research on Language Proficiency Assessment

### v1

@@ Appendix C Language-Specific Analysis
We provide in-depth analysis of model performances from the experiments in Section 4 across multiple dimensions of UniversalCEFR on results for select languages that we are qualified to interpret.
@@
English. Analysis of model performance shows that using fine-tuned models and linguistic feature-based classification (62% - 75%) obtains the best performance compared to prompting with instruction-tuned LLMs (19%-28%). However, these models tend to provide distinct patterns of specific CEFR labels. For the prompting setup, Gemma1, Gemma3, and EuroLLM models tend to give labels within the A1 and B1 range, while fine-tuned and feature-based models tend to lean towards the B1 and B2 range. For the pretrained and instruction-tuned models, this finding may be tied to A1 and B2 being the most common CEFR level band of most general-purpose texts found online, where the sources of the data from which these models are trained. For feature-based models, we note the potential effect of training and test data having higher instance counts for these level bands than A1, C1, and C2. Regarding model scale, upgraded versions from similar model families perform better than their previous versions, echoing previous findings in literature Imperial and Tayyar Madabushi (2024). This is particularly evident in Gemma3 being 12B in size and trained with massively multilingual data in 140+ languages and obtaining 28% in weighted F1 compared to Gemma1, which is 7B in size and English-centric, obtaining 21.8%. We note a potential default effect in using these models where additional specific CEFR descriptor information is not needed if the texts being evaluated are in English, due to the majority of data in the context of CEFR that is reflected in the training data being English.
@@
Spanish. Fine-tuned models outperform other setups, with feature-based approaches, especially Random Forest, achieving reasonable comparative performance. Moreover, multilingual models provide noticeable performance gains when compared to the English-only model. As per prompting strategy, for smaller multilingual models the language specific prompt seems to play a role in improving the performance as it also does for the Gemma1 English-only model, however the Gemma3 with 12B parameter is not affected by this and it’s been able to produce the best results of the LLMs (plus more sophisticated prompting strategies). As for the granularity of the input, models perform noticeably better at the document level than at the paragraph level, indicating that longer contexts are easier to classify than short ones. Finally, it is worth reporting a noticeable error of Gemma1: the prediction of C2 grade level, which does not exist in the Spanish dataset.
@@
Hindi. Both the Gemma models perform poorly compared to the fine-tuned XLM-R and the Random Forest variants and tend to classify most Hindi test items as A1 or A2. For example, Gemma1 puts 57% of Hindi test samples as A1, whereas there are only 19% of the test samples labeled as A1 in the gold standard labels. This is in line with the general trend noticed in Section 5.2, as the Hindi subset is entirely sentence-level. The distribution is closer to the Gold distribution for the fine-tuned and feature-engineered models. XLM-R fine-tuned models give the best performance amongst all models for Hindi, both in terms of exact category prediction and in terms of the degree of error (i.e., being within 1 level above or below the correct level). Finally, we looked at the correlation between a simple approximation of text length (calculated as the number of space-separated tokens), a commonly used variable in such automated language assessment approaches in NLP research, and the CEFR gold labels, as well as model-predicted labels, after converting them to a numeric scale. There was a high correlation between text length and the gold labels (0.7), which was also seen with the XLM-R model (0.74) and the Random Forest models (0.77). However, the Gemma models only had correlations of 0.44 and 0.54, respectively, with text length. However, considering that the Hindi subset only has sentence-level annotations without a larger context, it may be challenging to achieve further consistency with the gold standard labels, given the size of the annotated dataset. Future research should expand the available CEFR graded resources both in terms of quantity as well as granularity for the language.
@@
Portuguese. Comparing the different setups, we can see that the results for Portuguese follow the global tendency, with fine-tuned models achieving the highest performance, followed by feature-based models, and with prompting taking the last place. Although this study only covers paragraph-level learner data for Portuguese, similar patterns were observed on reference data Ribeiro et al. (2024b). However, comparing the results with those of other languages and, particularly, those with paragraph-level learner data, we can see that Portuguese is the language with the lowest performance ($\approx$33.5%). Several factors may contribute to this outcome. For instance, Portuguese is one of the languages with the least available training data, and the distribution of proficiency labels is right-skewed (especially in COPLE2). Furthermore, the data consists of texts written by learners from a wide range of L1 backgrounds with generally low proficiency. This makes it more difficult for models to identify consistent patterns due to strong L1 interference and low coverage. Overall, both fine-tuned and feature-based models seem to be unable to distinguish between sublevels, with most examples of both A levels being predicted as A1, and the remainder (mostly examples of the B levels) as B1. On the positive side, contrary to what was observed for other languages, the models do not seem to be influenced by text length, with the predictions of XML-R having a correlation of just 0.39 with that feature. The prompting approaches lead to a bias towards the prediction of levels A2 and B1, with the top performer among these approaches (Gemma3 with En-Write prompt) predicting A2 for 28% of the examples and B1 for 62%. Notably, when using the more descriptive prompts, the Gemma 1 model outperformed EuroLLM, in spite of having less parameters and not being specifically trained on Portuguese data.
@@
German. For German, the fine-tuned models (>70%) have been shown to outperform all other approaches, such as feature-based ($\approx$50%-65%) and prompting ($\approx$38%-46%), despite the presence of unbalanced CEFR levels in both the training and test data. The findings derived from the English-only and multilingual models, including fine-tuning and prompting methodologies, exhibit no notable difference. This may be due to the similarities between English and German, both of which are West Germanic languages. Alternatively, the great transferability of the fine-tuned English-only model may also be due to the large amount of German training data available (27,000 training samples). The feature-based models performed second best and were still able to compete with the fine-tuned models to some extent. This is surprising, given that a previous analysis showed that the features only exhibited low correlations with CEFR levels (see Section E.3). Proficiency assessment for German appears to require certain idiosyncratic features. For example, the feature covering the maximum distance between words in a dependency tree showed a high feature importance only for German, reflecting the language’s free word order and long-distance dependencies. For the prompting setup, the multilingual Gemma3 model performed, achieving good results for lower CEFR levels, but underpredicting higher levels. By contrast, Gemma1 significantly overpredicts level A1 (250 against 14 from the gold labels), resulting in poorer performance on average and across the other levels. One deceptive indicator might be the length of the texts to be classified, as reflected by the strong correlation between text length and Gemma1’s predictions ($r$=0.61). When comparing the prompting setups with regard to language-specific task descriptions, no clear trend emerges across all three LLMs, mirroring the difficulty of prompt engineering for a complex task such as multi-lingual proficiency classification.
@@ Appendix D Standardized Dataset Fields
Appendix E Full Linguistic Feature Information
@@
Table 16: List of linguistic features occurring in the top 10 of at least three languages. We use this list for the TopFeatures subset used in the experiment result in Table 3.
@@ E.2 Top Linguistic Features
To extract the top linguistic features (TopFeats), we selected those that are present in the top 10 ranked most important features for at least three languages. Using this criteria, we came up with a list of 23 linguistic features as reported in Table 16 which was then used in the experiment result in Table 3.
@@ Appendix F Hyperparameter Values
discourse ratio_referential Ratio of referential tokens to all tokens based on dependency tree relations SpaCy, Stanza, TSEval 0.385
@@
Length doc_num_sents Number of sentences per document / text SpaCy, Stanza, 0.3934 0.176
doc_num_tokens Number of tokens per document / text SpaCy, Stanza, -0.4254 0.2398
num_characters Number of characters per document / text SpaCy, Stanza, TSEval 0.1819 0.181
num_characters_per_sentence Number of characters per sentence SpaCy, Stanza, TSEval 0.541 0.2242
num_characters_per_word Number of characters per word SpaCy, Stanza, TSEval -0.3962 0.2414
num_sentences Number of sentences per document / text SpaCy, Stanza, TSEval 0.385
@@
num_syllables_in_sentence Number of syllables in document / text SpaCy, Stanza, pyphen, TSEval 0.219 0.1423
num_syllables_per_sentence Number of syllables per sentence SpaCy, Stanza, pyphen, TSEval -0.685 0.737
num_syllables_per_word Number of syllables per word SpaCy, Stanza, pyphen, TSEval 0.2394
@@
num_words Number of tokens per document / text SpaCy, Stanza, TSEval 0.658 0.298
num_words_per_sentence Number of tokens per sentence SpaCy, Stanza, TSEval
Lexical average_pos_in_freq_table Average frequency rank of tokens in FastText embeddings SpaCy, Stanza, TSEval
lexical_complexity_score Lexical complexity based on ranks of FastText embeddings SpaCy, Stanza, FastText, TSEval
type_token_ratio Type-token-Ratio SpaCy, Stanza, TSEval
max_pos_in_freq_table Maximum frequency rank of tokens in FastText embeddings SpaCy, Stanza, TSEval 0.1293
@@
Psycholinguistic concreteness Concreteness of words based on MEGAHR and FastText-Embeddings MEGA.HR crossling 0.1429 0.139
@@
imagebility imagebility of words based on MEGAHR and FastText-Embeddings MEGA.HR crossling 0.253 0.1635
Readability sentence_fkgl Flesch-Kincaid-Grading-Level, designed for English SpaCy, Stanza, TSEval 0.2459 0.1615
sentence_fre Flesch-Reading Ease, designed for English SpaCy, Stanza, TSEval 0.238 0.1572
@@
Syntactic avg_distance_betweeen_verb_particle Average distance between verb and particle based on dependency tree SpaCy, Stanza, -0.21 0.1543
@@
avg_distance_betweeen_words Average distance between words based on dependency tree SpaCy, Stanza, -0.59 0.179
@@
max_distance_betweeen_verb_particles Maximum distance between verb and particle based on dependency tree SpaCy, Stanza, -0.5 0.2121
@@
max_distance_betweeen_words Maximum distance between words based on dependency tree SpaCy, Stanza, 0.143 0.1557
@@
check_if_head_is_noun Whether the head of the dependency tree is a noun SpaCy, Stanza, TSEval -0.89 0.1385
@@
check_if_head_is_verb Whether the head of the dependency tree is a verb SpaCy, Stanza, TSEval -0.2226 0.1621
@@
check_if_one_child_of_root_is_subject Whether a child of a root is a subject (not a verb) SpaCy, Stanza, TSEval 0.2361 0.1428
@@
check_passive_voice Whether a sentence is in passive voice SpaCy, Stanza, TSEval 0.542 0.178
@@
average_length_NP Average length of noun phrase in tokens SpaCy, Stanza, TSEval -0.954 0.1394
@@
average_length_VP Average length of verb phrase in tokens SpaCy, Stanza, TSEval 0.1797 0.1353
@@
avg_length_PP Average length of prepositional phrase in tokens SpaCy, Stanza, TSEval 0.1278 0.854
@@
parse_tree_height Depth or height of the dependency tree SpaCy, Stanza, TSEval 0.3262 0.151
@@
ratio_clauses Ratio of tokens associated to a clause to all tokens based on dependency tree relations SpaCy, Stanza, TSEval -0.3364 0.1627
@@
ratio_of_coordinating_clauses Ratio of tokens associated to a coordinating clause to all tokens based on dependency tree relations SpaCy, Stanza, TSEval 0.883 0.1329
@@
ratio_of_subordinate_clauses Ratio of tokens associated to a subordinating clause to all tokens based on dependency tree relations SpaCy, Stanza, TSEval 0.214 0.746
@@
ratio_prepositional_phrases Ratio of tokens associated to a prepositional phrase to all tokens based on dependency tree relations SpaCy, Stanza, TSEval 0.4738 0.1873
@@
ratio_relative_phrases Ratio of tokens associated to a relative clause to all tokens based on dependency tree relations SpaCy, Stanza, TSEval -0.3976 0.1884
is_non_projective Whether a dependency tree is non projective SpaCy, Stanza, TSEval 0.51 0.253
Morphosyntactic ratio_Abbr_Yes Ratio of nouns which are an abbreviation to all nouns SpaCy, Stanza, UniversalDependencies 0.1822 0.869
ratio_Case_Abe Ratio of nouns in abessive case to all nouns SpaCy, Stanza, UniversalDependencies
ratio_Case_Acc Ratio of nouns in accusative case to all nouns SpaCy, Stanza, UniversalDependencies
ratio_Case_Ben Ratio of nouns in benefactive case to all nouns SpaCy, Stanza, UniversalDependencies
ratio_Case_Cau Ratio of nouns in causative case to all nouns SpaCy, Stanza, UniversalDependencies 0.1196
ratio_Case_Cmp Ratio of nouns in comparative case to all nouns SpaCy, Stanza, UniversalDependencies 0.3528 0.1248
ratio_Case_Cns Ratio of nouns in considerative case to all nouns SpaCy, Stanza, UniversalDependencies 0.224 0.224
ratio_Case_Com Ratio of nouns in comitative case to all nouns SpaCy, Stanza, UniversalDependencies -0.286 0.133
ratio_Case_Dat Ratio of nouns in dative case to all nouns SpaCy, Stanza, UniversalDependencies 0.2252
ratio_Case_Dis Ratio of nouns in distributive case to all nouns SpaCy, Stanza, UniversalDependencies
ratio_Case_Equ Ratio of nouns in equative case to all nouns SpaCy, Stanza, UniversalDependencies 0.5228
ratio_Case_Erg Ratio of nouns in ergative case to all nouns SpaCy, Stanza, UniversalDependencies
@@
Morphosyntactic ratio_Case_Par Ratio of nouns in partitive case to all nouns SpaCy, Stanza, UniversalDependencies
ratio_Case_Ins Ratio of nouns in instrumental case to all nouns (relevant for CZ) SpaCy, Stanza, UniversalDependencies 0.569
@@
ratio_Case_Ess Ratio of nouns in essive case to all nouns (relevant for ET) SpaCy, Stanza, UniversalDependencies
ratio_Case_Gen Ratio of nouns in genitive case to all nouns SpaCy, Stanza, UniversalDependencies
ratio_Case_Nom Ratio of nouns in nominative case to all nouns SpaCy, Stanza, UniversalDependencies -0.59
ratio_Case_Tem Ratio of nouns in temporal case to all nouns SpaCy, Stanza, UniversalDependencies 0.2172 0.119
@@
ratio_Case_Tra Ratio of nouns in translative case to all nouns (relevant for ET) SpaCy, Stanza, UniversalDependencies -0.243 0.145
@@
ratio_Case_Voc Ratio of nouns in vocative case to all nouns SpaCy, Stanza, UniversalDependencies -0.782 0.12
@@
ratio_Definite_Com Ratio of complex nouns to all nouns (relevant for AR) SpaCy, Stanza, UniversalDependencies 0.798
@@
ratio_Definite_Cons Ratio of nouns in construct state to all nouns (relevant for AR) SpaCy, Stanza, UniversalDependencies 0.115
ratio_Definite_Def Ratio of definite nouns to all nouns SpaCy, Stanza, UniversalDependencies 0.276 0.1269
ratio_Definite_Ind Ratio of indefinite nouns to all nouns SpaCy, Stanza, UniversalDependencies 0.811
ratio_Foreign_Yes Ratio of nouns which are foreign to all nouns SpaCy, Stanza, UniversalDependencies 0.2426 0.1892
@@
ratio_Mood_Cnd Ratio of verbs with conditional mood to all verbs SpaCy, Stanza, UniversalDependencies 0.16 0.1856
@@
ratio_Mood_Imp Ratio of verbs with imperative mood to all verbs SpaCy, Stanza, UniversalDependencies 0.2287 0.189
@@
ratio_Mood_Ind Ratio of verbs with indicative mood to all verbs SpaCy, Stanza, UniversalDependencies 0.2199
@@
ratio_Mood_Jus Ratio of verbs with jussive mood to all verbs SpaCy, Stanza, UniversalDependencies 0.223 0.727
@@
ratio_Mood_Qot Ratio of verbs with quotative mood to all verbs SpaCy, Stanza, UniversalDependencies 0.296 0.1296
@@
ratio_Mood_Sub Ratio of verbs with subjunctive mood to all verbs SpaCy, Stanza, UniversalDependencies 0.2782 0.1694
ratio_Number_Dual Ratio of nouns in dual number to all nouns SpaCy, Stanza, UniversalDependencies
@@
ratio_Number_Plur Ratio of nouns in plural number to all nouns SpaCy, Stanza, UniversalDependencies -0.1113 0.1636
@@
ratio_Number_Sing Ratio of nouns in singular number to all nouns SpaCy, Stanza, UniversalDependencies 0.225 0.1849
ratio_Polarity_Neg Ratio of negative verbs to all verbs SpaCy, Stanza, UniversalDependencies 0.682
ratio_Polarity_Pos Ratio of positive verbs to all verbs SpaCy, Stanza, UniversalDependencies 0.3196 0.178
ratio_Tense_Fut Ratio of verbs in future tense to all verbs SpaCy, Stanza, UniversalDependencies -0.97 0.2168
ratio_Tense_Imp Ratio of verbs in imperfect to all verbs SpaCy, Stanza, UniversalDependencies -0.1147 0.131
ratio_Tense_Past Ratio of verbs in past tense to all verbs SpaCy, Stanza, UniversalDependencies 0.954 0.1334
ratio_Tense_Pqp Ratio of verbs in pluferfect to all verbs SpaCy, Stanza, UniversalDependencies 0.144 0.1525
@@
ratio_Tense_Pres Ratio of verbs in present tenst to all verbs SpaCy, Stanza, UniversalDependencies
ratio_Voice_Act Ratio of verbs in active voice to all verbs SpaCy, Stanza, UniversalDependencies 0.5224 0.1874
ratio_Voice_Mid Ratio of verbs in middle voice to all verbs SpaCy, Stanza, UniversalDependencies 0.531 0.1656
@@
ratio_Voice_Pass Ratio of verbs in passive voice to all verbs SpaCy, Stanza, UniversalDependencies 0.3895 0.1613

### v2

@@ Appendix C Language-Specific Analysis
We provide in-depth analysis of model performances from the experiments in Section 5 across multiple dimensions of UniversalCEFR on results for selected languages that we are qualified to interpret.
@@
English. Analysis of model performance shows that using fine-tuned models and linguistic feature-based classification (62%-75%) obtains the best performance compared to prompting with instruction-tuned LLMs (19%-28%). However, these models tend to provide distinct patterns of specific CEFR labels. For the prompting setup, Gemma1, Gemma3, and EuroLLM models tend to give labels within the A1 and B1 range, while fine-tuned and feature-based models tend to lean towards the B1 and B2 range. For the pre-trained and instruction-tuned models, this finding may be tied to A1 and B2 being the most common CEFR level band of most general-purpose texts found online, where the sources of the data from which these models are trained. For feature-based models, we note the potential effect of training and test data having higher instance counts for these level bands than A1, C1, and C2. Regarding model scale, upgraded versions from similar model families perform better than their previous versions, echoing previous findings in literature Imperial and Tayyar Madabushi (2024). This is particularly evident in Gemma3 being 12B in size and trained with massively multilingual data in 140+ languages and obtaining 28% in weighted F1 compared to Gemma1, which is 7B in size and English-centric, obtaining 21.8%. We note a potential default effect in using these models where additional specific CEFR descriptor information is not needed if the texts being evaluated are in English, due to the majority of data in the context of CEFR that is reflected in the training data being English.
@@
Spanish. Fine-tuned models outperform other setups, with feature-based approaches, especially Random Forest, achieving reasonable comparative performance. Moreover, multilingual models provide noticeable performance gains when compared to the English-only model. As per prompting strategy, for smaller multilingual models the language-specific prompt seems to play a role in improving the performance as it also does for the Gemma1 English-only model, however, the Gemma3 with 12B parameter is not affected by this, and it has been able to produce the best results of the LLMs (plus more sophisticated prompting strategies). As for the granularity of the input, models perform noticeably better at the document level than at the paragraph level, indicating that longer contexts are easier to classify than short ones. Finally, it is worth reporting a noticeable error of Gemma1: the prediction of C2 grade level, which does not exist in the Spanish dataset.
@@
Hindi. Both the Gemma models perform poorly compared to the fine-tuned XLM-R and the Random Forest variants and tend to classify most Hindi test items as A1 or A2. For example, Gemma1 puts 57% of Hindi test samples as A1, whereas there are only 19% of the test samples labeled as A1 in the gold standard labels. This is in line with the general trend noticed in Section 6.2, as the Hindi subset is entirely sentence-level. The distribution is closer to the Gold distribution for the fine-tuned and feature-engineered models. XLM-R fine-tuned models give the best performance amongst all models for Hindi, both in terms of exact category prediction and in terms of the degree of error (i.e., being within 1 level above or below the correct level). Finally, we looked at the correlation between a simple approximation of text length (calculated as the number of space-separated tokens), a commonly used variable in such automated language assessment approaches in NLP research, and the CEFR gold labels, as well as model-predicted labels, after converting them to a numeric scale. There was a high correlation between text length and the gold labels (0.7), which was also seen with the XLM-R model (0.74) and the Random Forest models (0.77). However, the Gemma models only had correlations of 0.44 and 0.54, respectively, with text length. However, considering that the Hindi subset only has sentence-level annotations without a larger context, it may be challenging to achieve further consistency with the gold standard labels, given the size of the annotated dataset. Future research should expand the available CEFR-graded resources both in terms of quantity as well as granularity for the language.
@@
Portuguese. Comparing the different setups, we can see that the results for Portuguese follow the global tendency, with fine-tuned models achieving the highest performance, followed by feature-based models, and with prompting taking the last place. Although this study only covers paragraph-level learner data for Portuguese, similar patterns were observed on reference data Ribeiro et al. (2024b). However, comparing the results with those of other languages and, particularly, those with paragraph-level learner data, we can see that Portuguese is the language with the lowest performance ($\approx$33.5%). Several factors may contribute to this outcome. For instance, Portuguese is one of the languages with the least available training data, and the distribution of proficiency labels is right-skewed (especially in COPLE2). Furthermore, the data consists of texts written by learners from a wide range of L1 backgrounds with generally low proficiency. This makes it more difficult for models to identify consistent patterns due to strong L1 interference and low coverage. Overall, both fine-tuned and feature-based models seem to be unable to distinguish between sublevels, with most examples of both A levels being predicted as A1, and the remainder (mostly examples of the B levels) as B1. On the positive side, contrary to what was observed for other languages, the models do not seem to be influenced by text length, with the predictions of XML-R having a correlation of just 0.39 with that feature. The prompting approaches lead to a bias towards the prediction of levels A2 and B1, with the top performer among these approaches (Gemma3 with En-Write prompt) predicting A2 for 28% of the examples and B1 for 62%. Notably, when using the more descriptive prompts, the Gemma 1 model outperformed EuroLLM, in spite of having fewer parameters and not being specifically trained on Portuguese data.
@@
German. For German, the fine-tuned models (>70%) have been shown to outperform all other approaches, such as feature-based ($\approx$50%-65%) and prompting ($\approx$38%-46%), despite the presence of unbalanced CEFR levels in both the training and test data. The findings derived from the English-only and multilingual models, including fine-tuning and prompting methodologies, exhibit no notable difference. This may be due to the similarities between English and German, both of which are West Germanic languages. Alternatively, the great transferability of the fine-tuned English-only model may also be due to the large amount of German training data available (27,000 training samples). The feature-based models performed second best and were still able to compete with the fine-tuned models to some extent. This is surprising, given that a previous analysis showed that the features only exhibited low correlations with CEFR levels (see Section E.3). Proficiency assessment for German appears to require certain idiosyncratic features. For example, the feature covering the maximum distance between words in a dependency tree showed a high feature importance only for German, reflecting the language’s free word order and long-distance dependencies. For the prompting setup, the multilingual Gemma3 model performed, achieving good results for lower CEFR levels, but underpredicting higher levels. By contrast, Gemma1 significantly overpredicts level A1 (250 against 14 from the gold labels), resulting in poorer performance on average and across the other levels. One deceptive indicator might be the length of the texts to be classified, as reflected by the strong correlation between text length and Gemma1’s predictions (rr=0.61). When comparing the prompting setups with regard to language-specific task descriptions, no clear trend emerges across all three LLMs, mirroring the difficulty of prompt engineering for a complex task such as multi-lingual proficiency classification.
@@ Appendix D Standardized Dataset Fields
Appendix E Full Linguistic Feature Analysis
@@
Table 16: List of linguistic features occurring in the top 10 of at least three languages. We use this list for the TopFeatures subset used in the experiment result in Table 4.2.
@@ E.2 Top Linguistic Features
To extract the top linguistic features (TopFeats), we selected those that are present in the top 10 ranked most important features for at least three languages. Using this criteria, we came up with a list of 23 linguistic features as reported in Table 16 which was then used in the experiment result in Table 4.2.
@@ Appendix F Hyperparameter Values
[…]
@@
Length doc_num_sents Number of sentences per document / text SpaCy, Stanza, 0.1819 0.18
doc_num_tokens Number of tokens per document / text SpaCy, Stanza, 0.5041 0.22
num_characters Number of characters per document / text SpaCy, Stanza, TSEval 0.5224 0.19
num_characters_per_sentence Number of characters per sentence SpaCy, Stanza, TSEval 0.5301 0.17
num_characters_per_word Number of characters per word SpaCy, Stanza, TSEval 0.3895 0.16
num_sentences Number of sentences per document / text SpaCy, Stanza, TSEval -0.0324 0.11
@@
num_syllables_in_sentence Number of syllables in document / text SpaCy, Stanza, pyphen, TSEval 0.525 0.19
num_syllables_per_sentence Number of syllables per sentence SpaCy, Stanza, pyphen, TSEval 0.4634 0.23
num_syllables_per_word Number of syllables per word SpaCy, Stanza, pyphen, TSEval 0.3924 0.16
@@
num_words Number of tokens per document / text SpaCy, Stanza, TSEval 0.479 0.18
num_words_per_sentence Number of tokens per sentence SpaCy, Stanza, TSEval 0.4863 0.16
[…]
@@
Psycholinguistic concreteness Concreteness of words based on MEGAHR and FastText-Embeddings MEGA.HR crossling -0.4254 0.24
@@
imagebility imagebility of words based on MEGAHR and FastText-Embeddings MEGA.HR crossling -0.3962 0.24
[…]
@@
Syntactic avg_distance_betweeen_verb_particle Average distance between verb and particle based on dependency tree SpaCy, Stanza, 0.0385
@@
avg_distance_betweeen_words Average distance between words based on dependency tree SpaCy, Stanza, 0.3934 0.17
@@
max_distance_betweeen_verb_particles Maximum distance between verb and particle based on dependency tree SpaCy, Stanza, 0.0385
@@
max_distance_betweeen_words Maximum distance between words based on dependency tree SpaCy, Stanza, 0.2109 0.14
@@
check_if_head_is_noun Whether the head of the dependency tree is a noun SpaCy, Stanza, TSEval -0.1147 0.13
@@
check_if_head_is_verb Whether the head of the dependency tree is a verb SpaCy, Stanza, TSEval 0.0954 0.13
@@
check_if_one_child_of_root_is_subject Whether a child of a root is a subject (not a verb) SpaCy, Stanza, TSEval 0.0144 0.15
@@
check_passive_voice Whether a sentence is in passive voice SpaCy, Stanza, TSEval
@@
average_length_NP Average length of noun phrase in tokens SpaCy, Stanza, TSEval 0.3485 0.13
@@
average_length_VP Average length of verb phrase in tokens SpaCy, Stanza, TSEval 0.4324 0.16
@@
avg_length_PP Average length of prepositional phrase in tokens SpaCy, Stanza, TSEval 0.1406 0.07
@@
parse_tree_height Depth or height of the dependency tree SpaCy, Stanza, TSEval 0.4657 0.15
@@
ratio_clauses Ratio of tokens associated to a clause to all tokens based on dependency tree relations SpaCy, Stanza, TSEval 0.1918 0.12
@@
ratio_of_coordinating_clauses Ratio of tokens associated to a coordinating clause to all tokens based on dependency tree relations SpaCy, Stanza, TSEval 0.2503 0.16
@@
ratio_of_subordinate_clauses Ratio of tokens associated to a subordinating clause to all tokens based on dependency tree relations SpaCy, Stanza, TSEval 0.2361 0.14
@@
ratio_prepositional_phrases Ratio of tokens associated to a prepositional phrase to all tokens based on dependency tree relations SpaCy, Stanza, TSEval 0.1797 0.14
@@
ratio_relative_phrases Ratio of tokens associated to a relative clause to all tokens based on dependency tree relations SpaCy, Stanza, TSEval 0.3262 0.15
is_non_projective Whether a dependency tree is non projective SpaCy, Stanza, TSEval 0.0883 0.13
[…]
@@
Morphosyntactic ratio_Case_Ess Ratio of nouns in essive case to all nouns (relevant for ET) SpaCy, Stanza, UniversalDependencies 0.1196
ratio_Case_Gen Ratio of nouns in genitive case to all nouns SpaCy, Stanza, UniversalDependencies 0.3528 0.12
@@
ratio_Case_Ins Ratio of nouns in instrumental case to all nouns (relevant for CZ) SpaCy, Stanza, UniversalDependencies 0.2204 0.02
ratio_Case_Nom Ratio of nouns in nominative case to all nouns SpaCy, Stanza, UniversalDependencies -0.0286 0.13
ratio_Case_Par Ratio of nouns in partitive case to all nouns SpaCy, Stanza, UniversalDependencies 0.2252
ratio_Case_Tem Ratio of nouns in temporal case to all nouns SpaCy, Stanza, UniversalDependencies
@@
ratio_Case_Tra Ratio of nouns in translative case to all nouns (relevant for ET) SpaCy, Stanza, UniversalDependencies 0.5228
ratio_Case_Voc Ratio of nouns in vocative case to all nouns SpaCy, Stanza, UniversalDependencies
@@
ratio_Definite_Com Ratio of complex nouns to all nouns (relevant for AR) SpaCy, Stanza, UniversalDependencies
@@
ratio_Definite_Cons Ratio of nouns in construct state to all nouns (relevant for AR) SpaCy, Stanza, UniversalDependencies
ratio_Definite_Def Ratio of definite nouns to all nouns SpaCy, Stanza, UniversalDependencies 0.0569
ratio_Definite_Ind Ratio of indefinite nouns to all nouns SpaCy, Stanza, UniversalDependencies -0.059
ratio_Foreign_Yes Ratio of nouns which are foreign to all nouns SpaCy, Stanza, UniversalDependencies
@@
ratio_Mood_Cnd Ratio of verbs with conditional mood to all verbs SpaCy, Stanza, UniversalDependencies 0.2172 0.11
@@
ratio_Mood_Imp Ratio of verbs with imperative mood to all verbs SpaCy, Stanza, UniversalDependencies -0.0243 0.1
@@
ratio_Mood_Ind Ratio of verbs with indicative mood to all verbs SpaCy, Stanza, UniversalDependencies -0.0782 0.1
@@
ratio_Mood_Jus Ratio of verbs with jussive mood to all verbs SpaCy, Stanza, UniversalDependencies 0.0798
@@
ratio_Mood_Qot Ratio of verbs with quotative mood to all verbs SpaCy, Stanza, UniversalDependencies 0.1015
@@
ratio_Mood_Sub Ratio of verbs with subjunctive mood to all verbs SpaCy, Stanza, UniversalDependencies 0.2076 0.13
ratio_Number_Dual Ratio of nouns in dual number to all nouns SpaCy, Stanza, UniversalDependencies 0.0811
@@
ratio_Number_Plur Ratio of nouns in plural number to all nouns SpaCy, Stanza, UniversalDependencies 0.2426 0.19
@@
ratio_Number_Sing Ratio of nouns in singular number to all nouns SpaCy, Stanza, UniversalDependencies 0.0016 0.19
ratio_Polarity_Neg Ratio of negative verbs to all verbs SpaCy, Stanza, UniversalDependencies 0.2287 0.19
ratio_Polarity_Pos Ratio of positive verbs to all verbs SpaCy, Stanza, UniversalDependencies 0.2199
ratio_Tense_Fut Ratio of verbs in future tense to all verbs SpaCy, Stanza, UniversalDependencies 0.2023 0.07
ratio_Tense_Imp Ratio of verbs in imperfect to all verbs SpaCy, Stanza, UniversalDependencies 0.2096 0.13
ratio_Tense_Past Ratio of verbs in past tense to all verbs SpaCy, Stanza, UniversalDependencies 0.2782 0.17
ratio_Tense_Pqp Ratio of verbs in pluferfect to all verbs SpaCy, Stanza, UniversalDependencies
@@
ratio_Tense_Pres Ratio of verbs in present tenst to all verbs SpaCy, Stanza, UniversalDependencies -0.1113 0.16
ratio_Voice_Act Ratio of verbs in active voice to all verbs SpaCy, Stanza, UniversalDependencies 0.0225 0.18
ratio_Voice_Mid Ratio of verbs in middle voice to all verbs SpaCy, Stanza, UniversalDependencies 0.0682
@@
ratio_Voice_Pass Ratio of verbs in passive voice to all verbs SpaCy, Stanza, UniversalDependencies 0.3196 0.17